The dosage of a medicine ordered by a doctor is 40 mL of a 16% solution. A nurse has available both a 20% solution and a 4% solution of this medicine. How many milliliters of each could be mixed to prepare this 40-mL dosage?
Answer by Maths68(1474)

===============================================
Solution A
Amount = x
Concentration = 20% = 0.2
=============================================
Solution B
Amount = 40-x
Concentration = 4% = 0.04
====================================================
Resultant Solution
Amount = 40
Concentration = 16% = 0.16
========================================================
(Concentration of A)(Amount of A) + (Concentration of B)(Amount of B) = Concentration of Resultant)(Amount of Resultant)
(0.2)(x)+((0.04)(40-x))=(.16)(40)
0.16x+1.6=6.4
0.16x=6.4-1.6
0.16x=4.8
0.16x/0.16=4.8/0.16
x=30
===========================================
Solution A = 30ml
Solution B = 40-30=10ml
To get 40 ml of a 16% solution. 30 ml of 20% solution and a 10ml of 4% solution could be mixed